Risk taking is an important part of investing. In order to make suitable investment decisions on behalf of their customers, portfolio managers give a questionnaire to new customers to measure their desire to take financial risks. The scores on the questionnaire are approximately normally distributed with a mean of 50.5 and a standard deviation of 15. The customers with scores in the bottom 5% are described as "risk averse." What is the questionnaire score that separates customers who are considered risk averse from those who are not? Carry your intermediate computations to at least four decimal places. Round your answer to one decimal place.


Answer by Boreal(15235)

Need z-value for 5th percentile
that is -1.645
z=(x-mean)/sd
-1.64485=(x-50.5)/15
-24.675=(x-50.5)
x=25.827 when not rounded until the end.